# Groups of type $FP$ via graphical small cancellation

### 14 Dec 2023 v5

@@ 1 Introduction
This work was done while the first named author was working on his PhD under the supervision of the second named author. Further properties of the groups $G(S)$ and other methods for constructing spectacular complexes $K$ can be found in the PhD thesis of the first named author [9]. The main motivation for this work was the observation that the relations in the presentations for the groups $G_{L}(S)$ given in [23, Defn. 1.1] consist of a large family of ‘long’ relations together with a finite number of ‘short’ relations. Another motivation (which predates the work in [23]) was a conversation between the second named author and Martin Bridson, in which Martin Bridson pointed out that there ought to be other constructions of non-finitely presented groups of type $FP$ apart from that of Bestvina–Brady. The authors gratefully acknowledge this inspiration. The authors also gratefully acknowledge helpful comments on this work by Tim Riley. Finally, the authors thank the anonymous referee, whose numerous comments on three earlier versions have greatly improved the exposition and led to the correction of some minor errors.
@@ 2 Background
A graphical presentation is a 4-tuple $(\Gamma,L,\tau,\phi)$, where $\Gamma$ is a graph and $\phi$ is a reduced labelling of $\Gamma$ by $L$. The group $G(\Gamma,L,\tau,\phi)$ presented by a graphical presentation is the group given by the following ordinary presentation. The generators are the elements of $L$, subject to the following relations: for each $l\in L$, $\tau(l)$ is the inverse of $l$, and for every simple cycle in $\Gamma$, the word in $L$ obtained by going around that cycle is equal to the identity. This definition is used in [29, Theorem 1].
A reduced labelling of a graph $\Gamma$ by $L$ can be viewed as defining an immersion from $\Gamma$ to the rose $R_{L}$, which is the 1-vertex graph with directed edges in bijective correspondence with the elements of $L$. We remind the reader that $R_{L}$ has $|L|/2$ distinct 1-cells. The fundamental group of $R_{L}$ is of course a free group with representatives of the $\tau$-orbits in $L$ as free generators. The group $G(\Gamma,L,\tau,\phi)$ can be viewed as the quotient of the fundamental group of $R_{L}$ by the normal subgroup generated by the images of the fundamental groups of the components of $\Gamma$ under this immersion. This is the definition of $G(\Gamma,L,\tau,\phi)$ used in [18, Definition 1.1]. It is immediate that the two definitions describe the same group, essentially because the fundamental group of a graph can always be generated as a normal subgroup by simple cycles.
@@ 3 Using graphical small cancellation
There is a 3-dimensional $H(T)$-complex $F$ that contains both $E(S)/K_{S,T}$ and $E(T)$ as subcomplexes. The fundamental theorem of graphical small cancellation tells us that inside $E(T)$ there is a free $H(T)$-orbit of copies of cones on the degree $n$ subdivision of $K^{1}$, for each $n\in T-S$. These copies of the degree $n$ subdivision of $K^{1}$ are of course also present inside $E(S)/K_{S,T}$ since the graphical 1-skeleta are equal. However, in $E(S)/K_{S,T}$, instead of having a single cone vertex attached to each such copy, there is a specific subspace $K_{0}$ homeomorphic to $K$. We now describe $K_{0}$. The 1-skeleton of $K_{0}$ is the given copy of the degree $n$ subdivision of $K^{1}$. For each polygon $P$ of $K$, the degree $n$ subdivision of $\partial P$ is a relator in the given presentation for $H(S)$. Hence there is a 2-cell in $E(S)/_{S,T}$ whose boundary is the given copy of the degree $n$ subdivision of $\partial P$. These 2-cells (as $P$ ranges over all of the polygons in $K$) together with the 1-skeleton $K_{0}^{1}$ described earlier form the subspace $K_{0}$. To make $F$ from $E(S)/K_{S,T}$, attach a free orbit of cones on $K$ to these copies of $K$. The inclusion of $K^{1}$ into $K$ induces an embedding of the cone $C(K^{1})$ on $K^{1}$ into the cone $C(K)$, and hence an $H(T)$-equivariant inclusion $E(T)\rightarrow F$.
@@ 5 Embedding polygon subgroups
Let $K$ be a spectacular $2$-complex, let $Z={\mathbb{Z}}-\{0\}$ and let $L$ be the flag complex obtained from $K$ by replacing each polygon with the cone on its boundary. With the embedding $H_{P}\rightarrow G_{P}$ as described above there is an isomorphism, for each $S\subseteq{\mathbb{Z}}-\{0\}$, from $G(S)$ to the generalized Bestvina–Brady group $G_{L}(S\cup\{0\})$ in the sense of [23, Defn 1.1].
@@ 6 The projective line and its symmetries
A matrix whose characteristic polynomial is $(t-\lambda)(t-\mu)$ for $\lambda\neq\mu$ is conjugate to the diagonal matrix ${\mathrm{Diag}}(\lambda,\mu)$ via the change of basis that sends ${\bf e}_{1}$ to an eigenvector for $\lambda$ and ${\bf e}_{2}$ to an eigenvector for $\mu$. Let $g$ be the image of ${\mathrm{Diag}}(\lambda,\mu)$ inside $G$. Since $\lambda\neq\mu$, $g$ is not the identity element and so $g$ fixes just two points of ${\mathbb{P}}^{1}(q)$: the lines $\langle{\bf e}_{1}\rangle$ and $\langle{\bf e}_{2}\rangle$. Any element of $G$ that normalizes the subgroup generated by $g$ must either fix or exchange these two lines. The matrices that preserve the lines $\langle{\bf e}_{1}\rangle$ and $\langle{\bf e}_{2}\rangle$ are precisely the diagonal matrices, each of which centralizes ${\mathrm{Diag}}(\lambda,\mu)$, and the matrices that exchange these two lines are precisely the antidiagonal matrices. Each antidiagonal matrix conjugates ${\mathrm{Diag}}(\lambda,\mu)$ to ${\mathrm{Diag}}(\mu,\lambda)$. Since the product of ${\mathrm{Diag}}(\lambda,\mu)$ and ${\mathrm{Diag}}(\mu,\lambda)$ is a scalar matrix, any $h\in G$ that exchanges the two lines must conjugate $g$ to $g^{-1}$. Thus the normalizer in $G$ of the subgroup generated by $g$ is the image in $G$ of the group of diagonal and anti-diagonal matrices in $GL(2,q)$. This is a dihedral group of order $2(q-1)$. The cyclic subgroup of order $q-1$ fixes the same two points of ${\mathbb{P}}^{1}(q)$ as $g$ and centralizes $g$. The other elements of this dihedral group swap the two points fixed by $g$ and send $g$ to its inverse.
@@
An irreducible quadratic polynomial over ${\mathbb{F}}_{q}$ has both of its roots in the field ${\mathbb{F}}_{q^{2}}$ of $q^{2}$ elements. We may view $GL(2,q)$ as a subgroup of $GL(2,q^{2})$, and we may view $G$ as a subgroup of $PGL(2,q^{2})$. If $M$ is a matrix in $GL(2,q)$ whose characteristic polynomial is irreducible over ${\mathbb{F}}_{q}$, then over ${\mathbb{F}}_{q^{2}}$ we see that $M$ is conjugate to a diagonal matrix ${\mathrm{Diag}}(\lambda,\mu)$ for some $\mu\neq\lambda\in{\mathbb{F}}_{q^{2}}$. If $g$ is the image of this matrix in $PGL(2,q^{2})$, then we see that the normalizer in $PGL(2,q^{2})$ of this element is dihedral of order $2(q^{2}-1)$, and that the elements of this subgroup that fix the two points of ${\mathbb{P}}^{1}(q^{2})$ that are fixed by $g$ form a cyclic subgroup of order $q^{2}-1$, while the elements that swap these two points conjugate $g$ to $g^{-1}$. Apart from the two points fixed by $g$, the other points of ${\mathbb{P}}^{1}(q^{2})$ lie in orbits of length equal to the order of $g$. The action of $g$ on ${\mathbb{P}}^{1}(q^{2})$ preserves the subset ${\mathbb{P}}^{1}(q)\subseteq{\mathbb{P}}^{1}(q^{2})$. Since the two points of ${\mathbb{P}}^{1}(q^{2})$ that are fixed by $g$ are not contained ${\mathbb{P}}^{1}(q)$, the orbits for $g$ on ${\mathbb{P}}^{1}(q)$ are all of length equal to the order of $g$. Any matrix in $GL(2,q)$ that fixes the same two points of ${\mathbb{P}}^{1}(q^{2})$ as $g$ must also have an irreducible quadratic as its characteristic polynomial (since its eigenvectors in ${\mathbb{F}}_{q^{2}}^{2}$ do not lie in ${\mathbb{F}}_{q}^{2}$).
@@
So far we have an upper bound for the normalizer of a group element $g$ corresponding to a matrix $M$ whose characteristic polynomial is an irreducible quadratic, but we also need a lower bound. I.e., we need to construct a cyclic group of order $q+1$ that centralizes $g$ and a dihedral group of order $2(q+1)$ that normalizes the subgroup generated by $g$. Let $f(x)\in{\mathbb{F}}_{q}[x]$ be the characteristic polynomial of such an $M$. Define a (unital) ring homomorphism $\widetilde{\psi}_{M}:{\mathbb{F}}_{q}[x]\rightarrow M_{2}({\mathbb{F}}_{q})$ by $\widetilde{\psi}_{M}(x)=M$ and extending ${\mathbb{F}}_{q}$-linearly. By definition, the kernel of $\widetilde{\psi}_{M}$ is the ideal of ${\mathbb{F}}_{q}[x]$ generated by the minimal polynomial of $M$. Since the characteristic polynomial $f(x)$ is an irreducible quadratic, it is equal to the minimal polynomial of $M$ and the factor ring $E={\mathbb{F}}_{q}[x]/(f(x))$ is isomorphic to the field ${\mathbb{F}}_{q^{2}}$. Moreover, $\widetilde{\psi}_{M}$ induces an injective ${\mathbb{F}}_{q}$-linear homomorphism $\psi_{M}:E\rightarrow M_{2}({\mathbb{F}}_{q})$. Fix a non-zero vector ${\bf v}$ in ${\mathbb{F}}_{q}^{2}$ and define an ${\mathbb{F}}_{q}$-vector space homomorphism $\theta_{\bf v}:E\rightarrow{\mathbb{F}}_{q}^{2}$ by $\theta_{\bf v}(1)={\bf v}$ and $\theta_{\bf v}(x)=M{\bf v}$ where we have abused notation slightly by writing $x$ for the image of $x$ inside $E$. Since $E$ is being used as both a module and as a ring, we use the notation $E^{1}$ for the vector space and $M_{1}(E)$ as its endomorphism ring. The map $\theta_{\bf v}:E^{1}\rightarrow{\mathbb{F}}_{q}^{2}$ is an isomorphism of vector spaces over ${\mathbb{F}}_{q}$ and it is compatible with $\psi_{M}:M_{1}(E)\rightarrow M_{2}({\mathbb{F}}_{q})$ in the sense that for all $\lambda,\mu\in E$, $\theta_{\bf v}(\lambda\mu)=\psi_{M}(\lambda)\theta_{\bf v}(\mu)$. In particular, specializing to the case $\lambda=x$ gives that $\theta_{\bf v}(x\mu)=M\theta_{\bf v}(\mu)$ for any $\mu$, and so $\theta_{\bf v}x\theta_{\bf v}^{-1}=M$.
@@
The multiplicative group $GL(1,E)\leq M_{1}(E)$ is a cyclic group of order $q^{2}-1$ containing $x$. The Galois group of $E$ as an extension of ${\mathbb{F}}_{q}$ is a cyclic group of order two, generated by the Frobenius map $\phi_{q}:\lambda\mapsto\lambda^{q}$. This map is an ${\mathbb{F}}_{q}$-linear automorphism of $E$ of order two which normalizes $GL(1,E)$. Let $H$ be the group of ${\mathbb{F}}_{q}$-linear automorphisms of $E^{1}$ generated by $GL(1,E)$ and $\phi_{q}$. Thus $H$ is a group of ${\mathbb{F}}_{q}$-linear automorphisms of $E$ that has order $2(q^{2}-1)$, that normalizes the subgroup generated by $x$, and contains an index two cyclic subgroup that contains $x$. Now $\theta_{\bf v}:E^{1}\rightarrow{\mathbb{F}}_{q}^{2}$ is an isomorphism of ${\mathbb{F}}_{q}$-vector spaces such that $\theta_{\bf v}x\theta_{\bf v}^{-1}=M\in M_{2}({\mathbb{F}}_{q})$. Hence the group $H^{\prime}=\theta_{\bf v}H\theta_{\bf v}^{-1}\leq GL(2,q)$ normalizes the matrix $M$ and $M$ is contained in a cyclic subgroup of $H^{\prime}$ of order $q^{2}-1$. By the argument given earlier, the $(q+1)$st power of any element in this cyclic group is a scalar matrix, and so the image of $H^{\prime}$ in $G$ is a dihedral group of order $2(q+1)$, with $g$ (the image of $M$) contained in a cyclic subgroup of order $q+1$.

### 07 Feb 2024 v6

@@ 1 Introduction
This work was done while the first named author was working on his PhD under the supervision of the second named author. Further properties of the groups $G(S)$ and other methods for constructing spectacular complexes $K$ can be found in the PhD thesis of the first named author [9]. The main motivation for this work was the observation that the relations in the presentations for the groups $G_{L}(S)$ given in [23, Defn. 1.1] consist of a large family of ‘long’ relations together with a finite number of ‘short’ relations. Another motivation (which predates the work in [23]) was a conversation between the second named author and Martin Bridson, in which Martin Bridson pointed out that there ought to be other constructions of non-finitely presented groups of type $FP$ apart from that of Bestvina–Brady. The authors gratefully acknowledge this inspiration. The authors also gratefully acknowledge helpful comments on this work by Tim Riley. Finally, the authors thank the anonymous referee, whose numerous comments on four earlier versions have greatly improved the exposition and led to the correction of some minor errors.
@@ 2 Background
A graphical presentation is a 4-tuple $(\Gamma,L,\tau,\phi)$, where $\Gamma$ is a graph and $\phi$ is a reduced labelling of $\Gamma$ by $L$. The group $G(\Gamma,L,\tau,\phi)$ presented by a graphical presentation is the group given by the following ordinary presentation. The generators are the elements of $L$, subject to the following relations: for each $l\in L$, $\tau(l)$ is the inverse of $l$, and for every simple cycle in $\Gamma$, the word in $L$ obtained by going around that cycle is equal to the identity. This definition is used in [29, Thm. 1].
A reduced labelling of a graph $\Gamma$ by $L$ can be viewed as defining an immersion from $\Gamma$ to the rose $R_{L}$, which is the 1-vertex graph with directed edges in bijective correspondence with the elements of $L$. We remind the reader that $R_{L}$ has $|L|/2$ distinct 1-cells. The fundamental group of $R_{L}$ is of course a free group with representatives of the $\tau$-orbits in $L$ as free generators. The group $G(\Gamma,L,\tau,\phi)$ can be viewed as the quotient of the fundamental group of $R_{L}$ by the normal subgroup generated by the images of the fundamental groups of the components of $\Gamma$ under this immersion. This is the definition of $G(\Gamma,L,\tau,\phi)$ used in [18, Defn. 1.1]. It is immediate that the two definitions describe the same group, essentially because the fundamental group of a graph can always be generated as a normal subgroup by simple cycles.
@@ 3 Using graphical small cancellation
There is a 3-dimensional $H(T)$-complex $F$ that contains both $E(S)/K_{S,T}$ and $E(T)$ as subcomplexes. The fundamental theorem of graphical small cancellation tells us that inside $E(T)$ there is a free $H(T)$-orbit of copies of cones on the degree $n$ subdivision of $K^{1}$, for each $n\in T-S$. These copies of the degree $n$ subdivision of $K^{1}$ are of course also present inside $E(S)/K_{S,T}$ since the graphical 1-skeleta are equal. However, in $E(S)/K_{S,T}$, instead of having a single cone vertex attached to each such copy, there is a specific subcomplex $K_{0}$ homeomorphic to $K$, which has this copy of the degree $n$ subdivision of $K^{1}$ as its 1-skeleton, and in which each polygon $P$ of $K$ is attached to the degree $n$ subdivision of $\partial P$. To make $F$ from $E(S)/K_{S,T}$ attach a free $E(T)$-orbit of cones on $K$ to the $E(T)$-orbit of $K_{0}$. The inclusion of $K^{1}$ into $K$ induces an embedding of the cone $C(K^{1})$ on $K^{1}$ into the cone $C(K)$, and hence an $H(T)$-equivariant inclusion $E(T)\rightarrow F$.
@@ 5 Embedding polygon subgroups
Let $K$ be a spectacular $2$-complex, let $Z={\mathbb{Z}}-\{0\}$ and let $L$ be the flag complex obtained from $K$ by replacing each polygon with the cone on its boundary. With the embedding $H_{P}\rightarrow G_{P}$ as described above there is an isomorphism, for each $S\subseteq{\mathbb{Z}}-\{0\}$, from $G(S)$ to the generalized Bestvina–Brady group $G_{L}(S\cup\{0\})$ in the sense of [23, Defn. 1.1].
@@ 6 The projective line and its symmetries
A matrix whose characteristic polynomial is $(t-\lambda)(t-\mu)$ for $\lambda\neq\mu$ is conjugate to the diagonal matrix ${\mathrm{Diag}}(\lambda,\mu)$ via the change of basis that sends ${\bf e}_{1}$ to an eigenvector for $\lambda$ and ${\bf e}_{2}$ to an eigenvector for $\mu$. Let $g$ be the image of ${\mathrm{Diag}}(\lambda,\mu)$ inside $G$. Since $\lambda\neq\mu$, $g$ is not the identity element and so $g$ fixes just two points of ${\mathbb{P}}^{1}(q)$: the lines $\langle{\bf e}_{1}\rangle$ and $\langle{\bf e}_{2}\rangle$. Any element of $G$ that normalizes the subgroup generated by $g$ must either fix or exchange these two lines. The matrices that preserve the lines $\langle{\bf e}_{1}\rangle$ and $\langle{\bf e}_{2}\rangle$ are precisely the diagonal matrices, each of which centralizes ${\mathrm{Diag}}(\lambda,\mu)$, and the matrices that exchange these two lines are precisely the antidiagonal matrices. Each antidiagonal matrix conjugates ${\mathrm{Diag}}(\lambda,\mu)$ to ${\mathrm{Diag}}(\mu,\lambda)$. Since the product of ${\mathrm{Diag}}(\lambda,\mu)$ and ${\mathrm{Diag}}(\mu,\lambda)$ is a scalar matrix, any $h\in G$ that exchanges the two lines must conjugate $g$ to $g^{-1}$. Thus the normalizer in $G$ of the subgroup generated by $g$ is the image in $G$ of the group of diagonal and antidiagonal matrices in $GL(2,q)$. This is a dihedral group of order $2(q-1)$. The cyclic subgroup of order $q-1$ fixes the same two points of ${\mathbb{P}}^{1}(q)$ as $g$ and centralizes $g$. The other elements of this dihedral group swap the two points fixed by $g$ and send $g$ to its inverse.
@@
An irreducible quadratic polynomial over ${\mathbb{F}}_{q}$ has both of its roots in the field ${\mathbb{F}}_{q^{2}}$ of $q^{2}$ elements. We may view $GL(2,q)$ as a subgroup of $GL(2,q^{2})$, and we may view $G$ as a subgroup of $PGL(2,q^{2})$. If $M$ is a matrix in $GL(2,q)$ whose characteristic polynomial is irreducible over ${\mathbb{F}}_{q}$, then over ${\mathbb{F}}_{q^{2}}$ we see that $M$ is conjugate to a diagonal matrix ${\mathrm{Diag}}(\lambda,\mu)$ for some $\mu\neq\lambda\in{\mathbb{F}}_{q^{2}}$. The study of the properties of diagonalizable matrices in the previous two paragraphs therefore applies to $M$ as an element of $GL(2,q^{2})$. If $g$ is the image of the matrix $M$ in $PGL(2,q^{2})$, then we see that the normalizer in $PGL(2,q^{2})$ of this element is dihedral of order $2(q^{2}-1)$, and that the elements of this subgroup that fix the two points of ${\mathbb{P}}^{1}(q^{2})$ that are fixed by $g$ form a cyclic subgroup of order $q^{2}-1$, while the elements that swap these two points conjugate $g$ to $g^{-1}$. Arguing as in the previous paragraph, we see that a generator for this cyclic group acts as a single $(q^{2}-1)$-cycle on the complement of the two fixed points. Hence the points of ${\mathbb{P}}^{1}(q^{2})$ that are not fixed by $g$ lie in $g$-orbits of length equal to the order of $g$. The action of $g$ on ${\mathbb{P}}^{1}(q^{2})$ preserves the subset ${\mathbb{P}}^{1}(q)\subseteq{\mathbb{P}}^{1}(q^{2})$. Since the two points of ${\mathbb{P}}^{1}(q^{2})$ that are fixed by $g$ are not contained in ${\mathbb{P}}^{1}(q)$, the orbits for $g$ on ${\mathbb{P}}^{1}(q)$ are all of length equal to the order of $g$. Any matrix in $GL(2,q)$ that fixes the same two points of ${\mathbb{P}}^{1}(q^{2})$ as $g$ must also have an irreducible quadratic as its characteristic polynomial (since its eigenvectors in ${\mathbb{F}}_{q^{2}}^{2}$ do not lie in ${\mathbb{F}}_{q}^{2}$).
@@
So far we have an upper bound for the normalizer of a group element $g$ corresponding to a matrix $M$ whose characteristic polynomial is an irreducible quadratic, but we also need a lower bound. I.e., we need to construct a cyclic group of order $q+1$ that centralizes $g$ and a dihedral group of order $2(q+1)$ that normalizes the subgroup generated by $g$. Let $f(x)\in{\mathbb{F}}_{q}[x]$ be the characteristic polynomial of such an $M$. Define a (unital) ring homomorphism $\widetilde{\psi}_{M}:{\mathbb{F}}_{q}[x]\rightarrow M_{2}({\mathbb{F}}_{q})$ by $\widetilde{\psi}_{M}(x)=M$ and extending ${\mathbb{F}}_{q}$-linearly. By definition, the kernel of $\widetilde{\psi}_{M}$ is the ideal of ${\mathbb{F}}_{q}[x]$ generated by the minimal polynomial of $M$. Since the characteristic polynomial $f(x)$ is an irreducible quadratic, it is equal to the minimal polynomial of $M$ and the factor ring $E={\mathbb{F}}_{q}[x]/(f(x))$ is isomorphic to the field ${\mathbb{F}}_{q^{2}}$. Moreover, $\widetilde{\psi}_{M}$ induces an injective ${\mathbb{F}}_{q}$-linear homomorphism $\psi_{M}:E\rightarrow M_{2}({\mathbb{F}}_{q})$. Fix a non-zero vector ${\bf v}$ in ${\mathbb{F}}_{q}^{2}$ and define an ${\mathbb{F}}_{q}$-vector space homomorphism $\theta_{\bf v}:E\rightarrow{\mathbb{F}}_{q}^{2}$ by $\theta_{\bf v}(1)={\bf v}$ and $\theta_{\bf v}(x)=M{\bf v}$ where we have abused notation slightly by writing $x$ for the image of $x$ inside $E$. Since $E$ is being used as both a module and as a ring, we use the notation $E^{1}$ for $E$ viewed as a 1-dimensional vector space over $E$ and $M_{1}(E)$ for $E$ viewed as the endomorphism ring of $E^{1}$. Since $M$ has no eigenvalues in ${\mathbb{F}}_{q}$, ${\bf v}$ and $M{\bf v}$ are linearly independent and so the map $\theta_{\bf v}:E^{1}\rightarrow{\mathbb{F}}_{q}^{2}$ is an isomorphism of vector spaces over ${\mathbb{F}}_{q}$. Moreover it is compatible with $\psi_{M}:M_{1}(E)\rightarrow M_{2}({\mathbb{F}}_{q})$ in the sense that for all $\lambda,\mu\in E$, $\theta_{\bf v}(\lambda\mu)=\psi_{M}(\lambda)\theta_{\bf v}(\mu)$. In particular, specializing to the case $\lambda=x$ gives that $\theta_{\bf v}(x\mu)=M\theta_{\bf v}(\mu)$ for any $\mu$, and so $\theta_{\bf v}x\theta_{\bf v}^{-1}=M$. To avoid any confusion, we emphasize that this equation represents an identity between ${\mathbb{F}}_{q}$-linear automorphisms of ${\mathbb{F}}_{q}^{2}$, and that the symbol $x$ in this equation represents the map $\lambda\mapsto x\lambda$ for all $\lambda\in E^{1}$ which is $E$-linear and hence a fortiori ${\mathbb{F}}_{q}$-linear.
@@
The multiplicative group $GL(1,E)\leq M_{1}(E)$ is a cyclic group of order $q^{2}-1$ containing $x$. The Galois group of $E$ as an extension of ${\mathbb{F}}_{q}$ is a cyclic group of order two, generated by the Frobenius map $\phi_{q}:\lambda\mapsto\lambda^{q}$. This map is an ${\mathbb{F}}_{q}$-linear automorphism of $E$ of order two which is not $E$-linear, so it normalizes $GL(1,E)$ but is not contained in $GL(1,E)$. Since $\phi_{q}(\lambda\mu)=\lambda^{q}\mu^{q}$, it follows that conjugation by $\phi_{q}$ acts on $GL(1,E)$ as the group automorphism $\lambda\mapsto\lambda^{q}$. Let $H$ be the group of ${\mathbb{F}}_{q}$-linear automorphisms of $E^{1}$ generated by $GL(1,E)$ and $\phi_{q}$. Thus $H$ is a group of ${\mathbb{F}}_{q}$-linear automorphisms of $E$ that has order $2(q^{2}-1)$, that normalizes the subgroup generated by $x$, and contains an index two cyclic subgroup $GL(1,E)$ that contains $x$. Now $\theta_{\bf v}:E^{1}\rightarrow{\mathbb{F}}_{q}^{2}$ is an isomorphism of ${\mathbb{F}}_{q}$-vector spaces such that $\theta_{\bf v}x\theta_{\bf v}^{-1}=M\in M_{2}({\mathbb{F}}_{q})$. Hence the group $H^{\prime}=\theta_{\bf v}H\theta_{\bf v}^{-1}\leq GL(2,q)$ normalizes the subgroup generated by the matrix $M$ and $M$ is contained in a cyclic subgroup $C$ of $H^{\prime}$ of order $q^{2}-1$. Since $M$ is a power of a generator for $C$ we see that each generator of $C$ has the same fixed points in ${\mathbb{P}}^{1}(q^{2})$ as $M$. These points do not lie in ${\mathbb{P}}^{1}(q)$ and so each generator of $C$ is irreducible over ${\mathbb{F}}_{q}$. It follows by an argument given earlier that the $(q+1)$st power of any generator of $C$ is a scalar matrix and so lies in the kernel of the map $GL(2,q)\rightarrow G$. Hence the image of $C$ in $G$ is cyclic of order at most $q+1$. Since this kernel of the map $GL(2,q)\rightarrow G$ is cyclic of order $q-1$, the image of $C$ is cyclic of order exactly $q+1$ and the image of $H^{\prime}$ in $G$ is a dihedral group of order $2(q+1)$, with $g$ (the image of $M$) contained in a cyclic subgroup of order $q+1$ (the image of $C$).